Site soil sensor histories

Note that I am using the period data for CS616 output (in uSeconds) at sites where this is available.

GLand

2007-06-05 (First TOA5)

  • 23 SWC (cs616_wcr(x) and 23 Tsoil (soil_water_T(x)) headers.
  • soil_water_T(x) fields seem to be some kind of artifact - the values are scaled right, but they seem to be logging wrong (usually they are all the same).
  • There is a Tsoil_Avg column in TOA5 data that appears to represent surface soil temperature.
  • There may be some confusion between SWC_O3_37.5 and SWC_O3_52.5 sensors - check.

2008-08-21

  • 1 CS616 fixed

2009-02-20

  • Program change
  • Stopped reading the 1 extra CS616
  • cs616 headers changed to (x,y) format
  • Added HFP and TCAV measurements
  • Still no Tsoil measurements except SoilT_Avg

2009-07-15

  • 20 thermocouples were installed at the site
  • Headers are mux25t_signal_Avg(x)
  • The depths are approximately the same as the CS616s, except that there are no sensors at G3 37.5 and 52.5

2009-08

  • Fire at the site. CS616 and TCs went to NAN

2010-2-24

  • Thermocouple multiplexer turned on and TC measurements appear to be working again shortly after this.

2010-05-12

  • 20 Thermocouples and 18 new echo EC5 sensors installed in 4 pits.
  • Thermocouples buried at all 5 depths in each pit, but there are no 52.5cm echo probes in the 2 Open pits.
  • Note that the TC headers did not change from mux25t_signal_Avg(x), which is problematic.
  • There are also 4 "dummy" columns, which don't seem to be working.

2011-3-10

  • Echo probes rewired so that Grass1 and Open 1 have all 5 depths and Open2 and Grass2 have only 2.5 to 37.5cm depths.

2011-3-23

  • 20 CS616s were installed in the 4 pits alongside the Echo probes and TCs, but they are not read until June.

2011-06-09 to 2011-07-27

  • New programs to read the CS616 probes, and there was some downtime during this changeover.
  • Echo probes in Open pits were not read and headers disappeared after this change.
  • CS616s were briefly read into mux_soil_wcr(1-20) headers (3 days), then into swc_open1_depthincm_Avg.
  • Both of these headers persisted after July, but only the swc_... have data.

2012-08-09

  • Program change that reduced soil measurement cycling from 5min to 30min frequency.

2014-01-16

  • New program with new headers: SWC_G1_2p5_Avg, SWC_echo_G1_2p5_Avg, SoilT_G1_2p5_Avg

2014-01-17 and 2014-02-17 (2 TOA5 files)

  • Tsoil headers are messed up. There are no O3 and G3 pits and the headers are mixed up with the current ones, so these need to be renamed

JSav

2007-05-04 (First TOA5)

  • No SWC or SoilT data.

2007-08-31

  • New program - SWC and Tsoil headers present
    • 20 Tsoil sensors logging under soilT_Avg(x) (but there are 28 headers), VWC logging under Period(X) and VWC(X), but both are NANs
  • There was a gap in the data between 8/6 and 8/31

2007-09-24

  • Now the SWC sensors are working (logging to VWC(X)).

2008-02-25

  • SWC headers changed to cs616_wcr(x), but no data logged yet.
  • There are also soil_water_T(1) headers, but these also have no data, and may just be an artifact.

2008-03-31

  • Data is now logging into cs616_wcr(x) and soil_water_T(x) fields - data look similar but are scaled differently (not sure what the _T field is)
  • Actually I think, in this case the_T field is calculated VWC, while the cs616_wcr(x) field is the period.

2008-06-10

  • SHF plates installed

2008-12-18 to 2009-02-02

  • Soil sensors down for a while here

2009-07-08

  • There has been a changeover in soil data and probe locations
  • Tsoil headers changed to SoilT_J1_2.5_Avg
  • Unfortunately the period headers cs616_wcr(x), stayed the same, but now refer to different locations. We'll have to change the old ones to _old in the code
  • The calculated VWC headers changed to:SWC_J1_2.5

Summer 2011

  • Intermittent problems with SWC sensors (recording NAN).

2011-08-11

  • MUX replaced and seems to have fixed problem.

Spring 2012

  • More intermittent CS616 problems

2012-04-27

  • CS616s removed from CR5000 and placed on a CR1000 in the same box.
    • swc_j1_2.5_avg headers are calculated VWC, PA_uS_Avg(1-18) are the cs616 periods
    • These data come in on separate TOA5 files that are in the soil directory.
  • Headers disappear from CR5k TOA5 files at this point

2014-01-10

  • Header changes for CR5k
    • sap_Avg(x) to sap_J1a_Avg
    • open_north_Avg to shf_open_north_Avg
    • Added soil CO2 measurements

2014-01-17

  • Header changes affecting the CR1000 soil moisture datalogger
    • cs616 headers changed to SWC_O1_2p5 format

2014-02-25 to 2014-02-27

  • Reinstalled CS616s in new pits (see new headers in format SWC_J1_2p5_Avg, SoilT_J1_2.5_Avg)
  • All CS616s are on the CR5000 again
  • Soil heat flux plates reinstalled in 4 (?) pits (3 pits in data headers).
  • TCAVs added in similar pits as well.

2014-03-28

  • Dug up and reinstalled CS616 and Tsoil at 5, 10, 30 cm. Headers changed on 4/15 (possibly earlier)

SLand

2007-05-30 (first TOA5)

  • 21 soil_water_T(1) and 21 cs616_wcr(x) headers, but they don't seem to be functioning correctly.
  • Not sure what depth/location that 21st SWC and Tsoil reading is from

2007-07-05

  • cs616's are functioning, but sensors 12 and 13 are all NANs
  • soil_water_T sensors are not giving reasonable data and all sensors are identical.
  • There is a Tsoil_avgcolumn that looks correct.

2009-02-20

  • New program: 22 SWC sensor columns with soil_water_T_Avg(1) and cs616_wcr_Avg(x) headers
  • Now there are 6 heat flux plates

2009-03-31

  • At this point some rewiring was done on the AM16/32 Mux where the CS 616s are connected
  • It seems that there were a couple problems:
    • There are only 20 probes, but 22 are being output by the datalogger program
    • Pit O2 sensors are wired out of order at the MUX, so they have been logged in the wrong headers
    • Not sure what to do about this yet

2009-06-24

  • 20 thermocouples wired in to AM25t mux - reading into mux25t_signal_Avg(1) headers
  • Old thermocouple headers and data still present, but still garbage.

2012-08-09

Program changes: Tsoil now read into soilt_S2_52.5_Avg headers. CS616s now read every 30 minutes.

New GLand

2009-08-06 (first TOA5 file)

  • There are 20 CS616s and 20 Tsoil sensors with headers: cs616_wcr(20), SWC_g1_2.5 (calculated VWC - not used), and SoilT_g1_2.5_Avg

2010-02-24

  • Header changes for SWC to cs616_wcr(1,1), SWC_ (calculated values) headers disappeared.

2010-03-12

  • 10 of the CS616s stopped being logged - Only Grass1 and Open1 probes are now being measured
  • Order of measurements changed to G1 O1 G2 O2 for Tsoil

2010-03-26

  • New CS616s were added and wired in the G1 O1 G2 O2 order, but it doesn't look like they are being logged yet.

2010-05-05

  • Wiring continues to change, but it doesn't seem to affect the TOA5 files.
  • Supposedly TC order is G1, O2, O1 G2

2010-05-12

  • Further rewiring: Supposedly CS616 and TC order is now G1, G2, O1, O2.
  • Headers haven't changed

2010-05-28

  • New program that should now measure new CS616s and thermocouples in the order they are actually wired (see above).
  • It may be worth swapping headers in the processing code to account for mistaken probe identity over the last 2 months.

2011-06-23

  • New program for new TCAV added (installed on 5-24 - in open pit)

2014-01-17

  • Some header changes (fairly minor)

PJ

PJ is different - soil, sapflow, and CO2 data are logged by a separate CR23x at the site. Data from this logger is periodically downloaded and saved.

2007-11-09 (first TOA5)

  • No soil water or temperature measurements yet.

2008-06-19

  • New headers for Echo probes and Soil T arrive, but values are invalid.
  • 24 Echo probes: echo_wcr(x,y)
  • 25 Tsoil probes: soilT_Avg(x)

2008-08-12

Values finally valid for Echo probes and Tsoil on main logger (fixed by someone - in logbook).

  • I still don't know anything about the pit structure/depths for these SWC or Tsoil probes

2009-03-22

Thermocouples installed, not logging yet

2009-05-12

  • TDR probes installed and connected to CR23X
  • Soil CO2 probes wired in
  • Heatflux plate locations noted.
  • Soil measurements on the cr23x begin logging at this date. Data for 2009-2013 are in cr23x compilation files originally assembled by Laura Morillas. Find these in the "yearly_cr23x_compilations" direcory

2009-06-17

Echo, Tsoil headers removed from main datalogger (data was junk for the past while...)

2013-07-12

IRT sensors added to main datalogger

2014-01-01

cr23x data now comes in as monthly files in the "cr23x_files" directory

2014-01-10

New program to main datalogger includes some header changes (but no soil sensors)

2014-03-21

Sent new program to main datalogger

2014-03-26

Sent new program with added IRT measurements

PJ_girdle

PJ_girdle also has soil, sapflow, and CO2 data are logged by a separate CR23x at the site. Data from this logger is periodically downloaded and saved.

2009-02-23 (first TOA5)

  • No soil water or temperature measurements on main datalogger.

2009-04-28

  • TDR probes installed and connected to CR23X
  • Soil CO2 probes wired in
  • Heatflux plate locations noted.
  • Soil measurements on the cr23x begin logging at this date. Data for 2009-2013 are in cr23x compilation files originally assembled by Laura Morillas. Find these in the "yearly_cr23x_compilations" directory

2009-05-28 to 2009-06-09

Problems with cr23x data - many sensors not reading.

2013-07-03

NDVI and IRT sensors added to main datalogger

2014-01-01

cr23x data now comes in as monthly files in the "cr23x_files" directory

2014-01-10

New program to main datalogger includes some header changes (T, RH, flux vars, but no soil sensors)

2014-03-26

Sent new program to main datalogger adds a few measurements (Wood psychrometers?)

PPine

PPine soil data were logged by....

2006-08-30 (first TOA5)

  • No soil water or temperature measurements yet.

2006-11-18

  • 4 CS107 thermistors were added - presumably these are measuring the soil?

2009-05-22

3 new soil heat flux measurements added (shf_Avg(1-3))

2010-10-07

  • TCAV and Hukesflux shf installed

2013-12-28

New soil data headers appear and major new header changes begin

2014-01-03

Some minor rearranging of sensor order in output

2014-01-09

SWC and Tsoil (107) measurements disappear

2014-05-23

SWC and Tsoil (107) measurements reappear in a slightly different order

MCon

MCon soil data were logged by....

2006-08-31 (first TOA5)

  • No soil water or temperature measurements yet.

2006-12-04

  • 4 CS107 thermistors were added - presumably these are measuring the soil?

2009-06-01

3 new soil heat flux measurements added (shf_Avg(1-3))

2010-06-18

  • TCAV and echo probe installed (echo_h2o_Avg & TCAV_Avg)

2013-11-12

Site rebuilt after the fire. Now logging Tsoil, SWC, soil CO2 and new soil heat flux plates/TCAVs

2014

Some minor rearranging of sensor order in output, and minor header changes